OpenStudy (anonymous):

A circle has a diameter of 6 millimeters. What are the circumference and area of the circle?

OpenStudy (anonymous):

pie (3)^2

OpenStudy (anonymous):

circumference =\[\pi d\]

OpenStudy (anonymous):

kk thxs

OpenStudy (anonymous):

r=d/2 area= pi*r^2

OpenStudy (anonymous):

circumference= 2*pi*r